OpenStudy (anonymous):

Please help! This question deals with Newton's method to find the shortest possible distance

OpenStudy (anonymous):

Here is the question

OpenStudy (anonymous):

What I don't get the most is part A, how do I write the formula ?? The rest seems easy enough, but I can't proceed without doing A

OpenStudy (anonymous):

Do I use the distance formula \[D=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (anonymous):

For part a yes I think that you just need to replace x2 with the x coordinate of the rower and x1 with the x coordinate of the point on shore then do the same with the two y coordinates y2 is the rower's y and y1 is the y coordinate on shore which happens to be x^2

OpenStudy (anonymous):

For part A I've got \[D=\sqrt{(x-0.2)^2+(y-2.7)^2}\] since y = x^2 \[D=\sqrt{(x-0.2)^2+(x^2-2.7)^2}\]

OpenStudy (anonymous):

then for part B i've got \[f(x)=((x-0.2)^2+(x^2-2.7)^2)^{1/2}\] \[f'(x)=1/2((x-0.2)^2+(x^2-2.7)^2)^{-1/2}\times (2(x-0.2)\times (1)+ 2(x^2-2.7) \times(2x))\] \[f'(x)=1/2((x-0.2)^2+(x^2-2.7)^2)^{-1/2}\times (2(x-0.2)+ 4x(x^2-2.7))\] \[f'(x)=1/2((x-0.2)^2+(x^2-2.7)^2)^{-1/2}\times (2x-0.4+ 4x^3-10.8x)\] \[f'(x)=1/2((x-0.2)^2+(x^2-2.7)^2)^{-1/2}\times (4x^3-8.8x-0.4)\] \[f'(x)=\frac{ 4x^3-8.8x-0.4 }{ 2((x-0.2)^2+(x^2-2.7)^2)^{1/2} }\] \[f'(x)=\frac{ 4x^3-8.8x-0.4 }{ 2\sqrt{(x-0.2)^2+(x^2-2.7)^2} }\]

OpenStudy (anonymous):

aaaaannnnnnnnnnnnnd im stuck :/

OpenStudy (anonymous):

college math?

OpenStudy (anonymous):

senior high school ... :/

OpenStudy (anonymous):

@chingmachine Here it goes my proposal. I have reviewed calculations now. Any inconsistency found or doubt you have, ask

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

for part see what does the 0 -> mean??

OpenStudy (anonymous):

and we weren't taught the Newton's method in advance they simply just gave us this task so i'll attempt to understand it through practise :)) THANK YOU SO MUCH EVERYONE

OpenStudy (anonymous):

I order to find minimum distance you have to make the derivative of L(x) equal to zero

OpenStudy (anonymous):

Newton´s method is a very simple method to find zeros of any function that always work,except when there are no solutions or solution is a maximum or a minimum of the function

OpenStudy (anonymous):

for part (d), what does each table indicate?

OpenStudy (anonymous):

It is the application of Newton's method. Each "k" is an iteration that helps you find Xk+1 from the value of Xk, applying the formula I have given

OpenStudy (anonymous):

You apply the iteration until you find the function approaches zero with the desired level of accuracy

OpenStudy (anonymous):

ohhh I understand that now, thank you once again

OpenStudy (anonymous):

and the highlighted one is the answer because the smallest y-value(s) corresponds to the global minimum.. is this correct?

OpenStudy (anonymous):

That´s correct for the last table. And for the iterations (three tables) the green shaded cells indicate where the iteration can stop

OpenStudy (anonymous):

Becuase you do not see changes in Xk in first and second decimal places and the function approaches zero

OpenStudy (anonymous):

this question sure is long

OpenStudy (anonymous):

It is but is a very nice problem indeed

OpenStudy (anonymous):

everything makes sense now

OpenStudy (anonymous):

good to hear that

OpenStudy (anonymous):

you're a lifesaver! we weren't given solutions for this so my friends and I have been worrying over this question all day

OpenStudy (anonymous):

enjoy it then!